Suppose that a researcher studying the weight of female college athletes wants to predict the weights based on height, measured in inches, the percentage of body fat of an athlete, and age. The researcher calculates the regression equation as (weight) = 4.362*(height) + 1.002*(percent body fat) - 1.187*(age) - 94.058. If a female athlete is 61 inches tall, has a 23 percentage of body fat, is 24 years old, and has a weight of 198.339. What is the residual?

Question 16 options:

1)

-31.757

2)

We do not know the observations in the data set, so we cannot answer that question.

3)

31.757

4)

-62.301

5)

105.582
Answer #1

weight = 4.362*(height) + 1.002*(percent body fat) - 1.187*(age) - 94.058

           = 4.362 * 61 + 1.002 * 23 - 1.187 * 24 - 94.058

           = 166.582

residual value = observed value - predicted value = 198.339 - 166.582 = 31.757

Option-3) 31.757